Monopoly Market: MC AC 4 2 MR 100 125 150 175 200 300 a. What is the profit maximizing output and price for this monopoly market? b. What is the monopoly profit? C. What would be the price and quantity if this was a perfectly competitive market? d. What is the deadweight loss, measured in dollars?
Answer #1

a) Profit is maximized at the level of output at which MR = MC. This gives Q* = 100 units. Price at 100 units is $9. This is the profit maximizing price and quantity for monopolist

b) Profit = (Price - ATC)*Q = (9 - 6)*100 = $300

c) For competitive firm, P = MC gives a price of $7 per unit and an output level of 175 units

d) DWL = 0.5*(Current price - MR at current quantity)*(competitive quantity - monopoly quantity)

= 0.5*(9 - 4)*(175 - 100)

= $187.50.